# thanks: These authors contributed equallythanks: These authors contributed equally

### v1

@@ Abstract
Negative energetic elasticity in gels challenges the conventional understanding of gel elasticity, and despite extensive research, a concise explanation remains elusive. In this Letter, we use the weakly self-avoiding walk (the Domb–Joyce model; DJ model) and interacting self-avoiding walk (ISAW) to investigate its emergence in polymer chains. Using exact enumeration, we show that both the DJ model and ISAW exhibit negative energetic elasticity, which is caused by effective soft-repulsive interactions between polymer segments. Moreover, we find that a universal scaling law for the internal energy of both models, with a common exponent of $7/4$, holds consistently across both random-walk–self-avoiding-walk and self-avoiding-walk–neighbor-avoiding-walk crossovers. These findings suggest that negative energetic elasticity is a fundamental and universal property of polymer networks and chains.
Gels are soft solids containing large amounts of solvent. They are used in the production of everyday items such as jellies and soft contact lenses; for many applications, a particular elastic modulus is required. Conventionally, the shear modulus $G$ of a gel is predicted using a classical rubber elasticity theory, such as the affine network [1], phantom network [2], or modified phantom network [3] model. These models assume that $G$ is predominantly determined by entropic elasticity and therefore that $G$ is approximately proportional to the absolute temperature ($T$) as $G\approx aT$. However, recent studies [4, 5] have discovered cases of negative energetic elasticity, in which $G=aT-b$ with a significantly large negative constant term $-b$ in some narrow range of temperature. Initially observed in hydrogels [4, 6, 7], negative energetic elasticity was subsequently confirmed in silicone gels [8] and calcium carbonate-based gels [9]. These findings have revealed a fundamental distinction between gels and rubbers: gels possess an intrinsic solvent-induced negative energetic contribution to the elasticity, whereas rubbers, lacking sufficient solvent content, do not. Thus, novel elastic theories beyond conventional paradigms aligning gel and rubber elasticity are needed.
@@
Figure 1: Configurations of 20-step (a) random walk (RW) $\omega_{\mathrm{a}}$, (b) self-avoiding walk (SAW) $\omega_{\mathrm{b}}$, and (c) neighbor-avoiding walk (NAW) $\omega_{\mathrm{c}}$ on cubic lattices with endpoints anchored at the origin and $(10,0,0)$. The red concentric circles in $\omega_{\mathrm{a}}$ indicate a site occupied by two segments, and the red oval in $\omega_{\mathrm{b}}$ highlights a pair of nearest-neighbor segments. (d) Inclusion relations $\Omega_{\mathrm{RW}}\supset\Omega_{\mathrm{SAW}}\supset\Omega_{\mathrm{NAW}}$, where $\Omega_{\mathrm{RW}}$, $\Omega_{\mathrm{SAW}}$, and $\Omega_{\mathrm{NAW}}$ are the configuration spaces of RWs, SAWs, and NAWs, respectively.
@@
In this Letter, we investigate the microscopic origin of negative energetic elasticity by analyzing the Domb–Joyce (DJ) model [14, 15, 16], i.e., a simple extension of the random walk (RW) [Fig. 1(a)] with short-range repulsions introduced at the intersections along the polymer chain. As de Gennes has noted [17], the RW is “one of the simplest idealizations of a flexible polymer chain” and serves as a starting point for understanding polymer behavior. The DJ model reduces to the self-avoiding walk (SAW) [18, 19] [Fig. 1(b)] in the zero-temperature limit; therefore, it can be viewed as a bridge between the RW and SAW, and it has been used to explore the differences in their critical exponents. Recent studies have employed variations of the DJ model (extending from a self-interacting chain to a time series of a non-Markovian walker) to describe agents such as cells, animals, and active matter [20, 21, 22, 23, 24].
@@
Through exact enumeration, we demonstrate that the DJ model exhibits negative energetic elasticity and provides a concise and intuitive interpretation of its origin in terms of the competition between the energetic and entropic contributions to the elastic response. We compare the DJ model with the interacting self-avoiding walk (ISAW) [18, 25] used in Ref. [10] and suggest that any polymer chain possessing effective self-repulsive interactions between its segments exhibits negative energetic elasticity. Furthermore, we find a universal scaling law for the internal energy of both the DJ model and ISAW, with a common exponent of $7/4$. These findings suggest that negative energetic elasticity is a universal property of polymeric chains and networks, arising from the interplay between the chain’s conformational entropy and the effective soft-repulsive interactions between its segments.
@@
Figure 2: Interaction energy ($\varDelta E$) in five lattice polymer models. The range of interactions is denoted by $d$; when $d=0$ or $1$, the interaction is between two segments occupying the same site or nearest-neighbor sites, respectively. The RW has no interactions between segments. The SAW introduces a hard-core repulsion ($d=0$) that prohibits segment overlap. The NAW introduces an additional hard-shell repulsion ($d=1$) between nearest-neighbor segments. The DJ model and the ISAW incorporate soft-core ($d=0$) and soft-shell ($d=1$) repulsions, respectively, with interaction strength $\varepsilon\,(>0)$. Varying $\varepsilon$ in these models allows continuous crossovers between the RW, SAW, and NAW.
@@
Models: RW–SAW and SAW–NAW crossovers.—We consider an $n$-step random walk (RW) on a simple cubic lattice with fixed endpoints $\omega(0)=(0,0,0)$ and $\omega(n)=(r,0,0)$, where $r$ is the end-to-end distance ($r\geq 0$) and the lattice spacing is set to $1$. The RW is defined by a sequence of sites $\omega=(\omega(0),\omega(1),\ldots,\omega(n))$ satisfying $\omega(i)\in\mathbb{Z}^{3}$ and $|\omega(i+1)-\omega(i)|=1$ for $i=0,1,\ldots,n-1$. A self-avoiding walk (SAW) is defined as an RW that satisfies $\omega(i)\neq\omega(j)$ for all $i\neq j$ [19]. Figures 1(a) and 1(b) show configurations of a $20$-step RW ($\omega_{\mathrm{a}}$) and SAW ($\omega_{\mathrm{b}}$), respectively.
@@
To investigate the effects of short-range repulsive interactions on the elasticity of the polymer chains, we introduce the DJ model and ISAW [25] as extensions of the RW and SAW, respectively. Both models use a unified energy function for the repulsive interaction:
@@
Figure 2 illustrates the interactions in five lattice polymer models, revealing the RW–SAW and SAW–NAW crossovers. The RW [Fig. 1(a)] has no interaction between its segments (top left of Fig. 2). The SAW [Fig. 1(b)] emerges from the RW when a hard-core repulsion is introduced (top center of Fig. 2) that prohibits segment overlap. The neighbor-avoiding walk (NAW) [Fig. 1(c)] emerges from the SAW when an additional hard-shell repulsion is introduced (top right of Fig. 2[26]. The RW, SAW, and NAW all exhibit purely entropic elasticity; energetic elasticity is excluded because all configurations have identical energies. By contrast, the DJ model and ISAW introduce energy gradients via soft-core and soft-shell repulsions, respectively, with interaction energies $\varepsilon\,(>0)$ (bottom panels of Fig. 2). By varying $\varepsilon$, the DJ model (ISAW) connects the RW (SAW) at $\varepsilon=0$ to the SAW (NAW) at $\varepsilon=\infty$. These energy gradients enable the DJ model and ISAW to exhibit both energetic and entropic elasticities.
Figure 3: Emergence of negative energetic elasticity in RW–SAW and SAW–NAW crossovers with $(n,r)=(20,10)$. The DJ model bridges the RW and SAW (left column), and the ISAW bridges the SAW and NAW (right column). The data for the ISAW are obtained from Ref. [10]. The top panels show the dependence of total stiffness $\beta k$ and the entropic contribution to stiffness $\beta k_{S}$ on interaction strength $\beta\varepsilon$. The middle panels show the energetic contribution to stiffness $\beta k_{U}$. The bottom panels show the ratio $-k_{U}/k$ and its first-order approximation $\tau\beta\varepsilon$ around $\beta\varepsilon=0$. The maximum of $-k_{U}/k$ occurs around $\beta\varepsilon\simeq 1$ for both models.
@@
where $\beta\,[\equiv 1/(k_{\mathrm{B}}T)]$ is the inverse temperature, $k_{\mathrm{B}}$ is the Boltzmann constant, $W_{n,m}(r)$ is the number of possible configurations $\omega$ for given $(n,r,m)$, and $m_{\mathrm{ub}}$ is an upper bound on $m$. The free energy, internal energy, and entropy are given by $A(r,\beta\varepsilon)=-(1/\beta)\ln Z(r,\beta\varepsilon)$, $U(r,\beta\varepsilon)=\varepsilon\,\sum_{m=0}^{m_{\mathrm{ub}}}mW_{n,m}(r)\,e^{-\beta\varepsilon m}$, and $S(r,\beta\varepsilon)=[A(r,\beta\varepsilon)-U(r,\beta\varepsilon)]/T$, respectively.
@@
where $\varDelta W_{n,m}(r)\equiv\left[W_{n,m}(r+\varDelta r)-W_{n,m}(r-\varDelta r)\right]/2$ and $\varDelta^{2}W_{n,m}(r)\equiv W_{n,m}(r+\varDelta r)-2W_{n,m}(r)+W_{n,m}(r-\varDelta r)$. Here, $\varDelta r\equiv 2$ because $\omega$ exists only when $r$ and $n$ are both even or both odd. The stiffness $k$ is the sum of energetic ($k_{U}$) and entropic ($k_{S}$) contributions, where $k_{U}(r,\beta\varepsilon)\equiv\partial^{2}U(r,\beta\varepsilon)/\partial r^{2}$ and $k_{S}(r,\beta\varepsilon)\equiv-T\partial^{2}S(r,\beta\varepsilon)/\partial r^{2}$. We use $k_{S}(r,\beta\varepsilon)=-\beta\varepsilon\,\partial k(r,\beta\varepsilon)/\partial(\beta\varepsilon)$, which is derived from Maxwell’s relations [1, 4, 5, 10], and $k_{U}=k-k_{S}$. When it is necessary to distinguish the models, we employ “RW” or “SAW” as superscripts on $W_{n,m}(r)$, and “DJ” and “ISAW” on other physical quantities.
@@
Exact enumeration for the DJ model and ISAW.—We exactly enumerated $W^{\mathrm{RW}}_{n,m}(r)$ for $n=1,\dots,20$, $W^{\mathrm{RW}}_{n,m}(n-8)$ for $n=21,22,23$, $W^{\mathrm{RW}}_{n,m}(n-10)$ for $n=21,\dots,26$, and $W^{\mathrm{SAW}}_{n,m}(n-10)$ for $n=21,\dots,29$, using the simplest recursive algorithm [27] with two pruning algorithms. Here, we considered the octahedral symmetry of the simple cubic lattice and the reachability of $\omega$ to an endpoint on the $x$-axis (the method is explained in Sec. S2 of Ref. [10]). These results are consistent with those reported in Refs. [10, 28, 29, 30, 31, 32, 33, 34].

### v2

@@ Abstract
Negative energetic elasticity in gels challenges the conventional understanding of gel elasticity; despite extensive research, a concise explanation remains elusive. In this study, we use the weakly self-avoiding walk (the Domb–Joyce model; DJ model) and interacting self-avoiding walk (ISAW) to investigate the emergence of negative energetic elasticity in polymer chains. Using exact enumeration, we show that both the DJ model and ISAW exhibit negative energetic elasticity, which is caused by effective soft-repulsive interactions between polymer segments. Moreover, we find that a universal scaling law for the internal energy of both models, with a common exponent of $7/4$, holds consistently across both random-walk–self-avoiding-walk and self-avoiding-walk–neighbor-avoiding-walk crossovers. These findings suggest that negative energetic elasticity is a fundamental and universal property of polymer networks and chains.
I Introduction
Polymer gels are soft solids containing large amounts of solvent. They are used in the production of everyday items such as jellies and soft contact lenses; for many applications, a particular elastic modulus is required. Conventionally, the shear modulus $G$ of a gel is predicted using a classical rubber elasticity theory, such as the affine network [1], phantom network [2], or modified phantom network [3] model. These models assume that $G$ is predominantly determined by entropic elasticity and therefore is approximately proportional to the absolute temperature ($T$), i.e., $G\approx aT$. However, recent studies [4, 5] discovered cases of negative energetic elasticity, in which G=aTbG=aT-b with a significantly large negative constant term b-b in some narrow range of temperature. Initially observed in hydrogels [4, 6, 7], negative energetic elasticity was subsequently confirmed in silicone gels [8] and calcium carbonate-based gels [9]. These findings revealed a fundamental distinction between gels and rubbers: gels possess an intrinsic solvent-induced negative energetic contribution to the elasticity, whereas rubbers, lacking sufficient solvent content, do not. Thus, it is necessary to develop novel elastic theories that go beyond the conventional paradigms aligning gel and rubber elasticity.
@@ I Introduction
Figure 1: Configurations of 20-step (a) RW $\omega_{\mathrm{a}}$, (b) SAW $\omega_{\mathrm{b}}$, and (c) NAW $\omega_{\mathrm{c}}$ on cubic lattices with endpoints anchored at the origin and $(10,0,0)$. The red concentric circles in $\omega_{\mathrm{a}}$ indicate a site occupied by two segments, and the red oval in $\omega_{\mathrm{b}}$ highlights a pair of nearest-neighbor segments. (d) Inclusion relations $\Omega_{\mathrm{RW}}\supset\Omega_{\mathrm{SAW}}\supset\Omega_{\mathrm{NAW}}$, where $\Omega_{\mathrm{RW}}$, $\Omega_{\mathrm{SAW}}$, and $\Omega_{\mathrm{NAW}}$ are the configuration spaces of RWs, SAWs, and NAWs, respectively.
@@
In this study, we investigate the microscopic origin of negative energetic elasticity by analyzing the Domb–Joyce (DJ) model [14, 15, 16], a simple extension of the random walk (RW) [Fig. 1(a)] with short-range repulsions introduced at the intersections along the polymer chain. As de Gennes noted [17], the RW is “one of the simplest idealizations of a flexible polymer chain” and serves as a starting point for understanding polymer behavior. The DJ model reduces to the self-avoiding walk (SAW) [18, 19] [Fig. 1(b)] in the zero-temperature limit; therefore, it can be viewed as a bridge between the RW and SAW, and it has been used to explore the differences in their critical exponents. Recent studies employed variations of the DJ model (extending from a self-interacting chain to a time series of a non-Markovian walker) to describe agents such as cells, animals, and active matter [20, 21, 22, 23, 24].
@@
Through exact enumeration, we demonstrate that the DJ model exhibits negative energetic elasticity and provides a concise and intuitive explanation of its origin in terms of the competition between the energetic and entropic contributions to the elastic response. We compare the DJ model with the interacting self-avoiding walk (ISAW) [18, 25] used in Ref. [10] and suggest that any polymer chain possessing effective self-repulsive interactions between its segments exhibits negative energetic elasticity. Furthermore, we find a universal scaling law for the internal energy of both the DJ model and ISAW, with a common exponent of $7/4$. These findings suggest that negative energetic elasticity is a universal property of polymer chains and networks, arising from the interplay between the chain’s conformational entropy and the effective soft-repulsive interactions between its segments.
The rest of this paper is organized as follows. In Sec. II, we define and relate five lattice polymer models (RW, SAW, NAW, DJ model, and ISAW) and present the framework to analyze their elastic properties using the finite-difference form of the stiffness. In Sec. III, we describe our exact enumeration methods for the DJ model and ISAW, validating our results through comparison with existing literature. In Sec. IV, we present polynomials that exactly reproduce the enumeration results for arbitrary chain lengths, enabling analytic calculations of physical quantities. In Sec. V, we demonstrate the emergence of negative energetic elasticity in RW–SAW and SAW–NAW crossovers and offer an intuitive interpretation based on entropic and energetic competition. In Sec. VI, we reveal a universal scaling law for the internal energy with an exponent of $7/4$ under the on-axis constraint. In Sec. VII, we summarize our findings and their implications for understanding the elasticity of polymer networks and chains. Appendixes A to C derive the upper bound of summations and present detailed polynomial expressions, and Appendix D extends our on-axis scaling analysis to off-axis end-to-end vectors. All numerical results in this paper are reproducible using the numbers provided in the tables in the Supplemental Material [27] and Appendix C together with those in the Supplemental Material of Ref. [10].
@@
Figure 2: Interaction energy ($\varDelta E$) in five lattice polymer models. The range of interactions is denoted by $d$; for $d=0$ or $1$, the interaction is between two segments occupying the same site or nearest-neighbor sites, respectively. The RW has no interactions between segments. The SAW introduces a hard-core repulsion ($d=0$) that prohibits segment overlap. The NAW introduces an additional hard-shell repulsion ($d=1$) between nearest-neighbor segments. The DJ model and the ISAW incorporate soft-core ($d=0$) and soft-shell ($d=1$) repulsions, respectively, with interaction strength $\varepsilon\,(>0)$. Varying $\varepsilon$ in these models allows continuous crossovers between the RW, SAW, and NAW.
II Models: RW–SAW and SAW–NAW crossovers
@@
We consider an $n$-step random walk (RW) on a simple cubic lattice with fixed endpoints $\omega(0)=(0,0,0)$ and $\omega(n)=(r_{x},r_{y},r_{z})\in\mathbb{Z}^{3}$. We define the end-to-end distance $r=|\omega(n)-\omega(0)|=\sqrt{r_{x}^{2}+r_{y}^{2}+r_{z}^{2}}$. We mainly focus on the on-axis constraint $\omega(n)=(r,0,0)$ (i.e., $r_{y}=r_{z}=0$) as a representative case. The lattice spacing is set to $1$. The RW is defined by a sequence of sites $\omega=[\omega(0),\omega(1),\ldots,\omega(n)]$ satisfying $\omega(i)\in\mathbb{Z}^{3}$ and $|\omega(i+1)-\omega(i)|=1$ for $i=0,1,\ldots,n-1$. A self-avoiding walk (SAW) is defined as an RW that satisfies $\omega(i)\neq\omega(j)$ for all $i\neq j$ [19]. Figures 1(a) and 1(b) show configurations of a $20$-step RW ($\omega_{\mathrm{a}}$) and SAW ($\omega_{\mathrm{b}}$), respectively.
@@
To investigate the effects of short-range repulsive interactions on the elasticity of the polymer chains, we introduce the DJ model and ISAW as extensions of the RW and SAW, respectively. Both models use a unified energy function for the repulsive interaction
@@
Figure 2 illustrates the interactions in five lattice polymer models, revealing the RW–SAW and SAW–NAW crossovers. The RW [Fig. 1(a)] has no interaction between its segments (top left of Fig. 2). The SAW [Fig. 1(b)] emerges from the RW when a hard-core repulsion is introduced (top center of Fig. 2) that prohibits segment overlap. The NAW [Fig. 1(c)] emerges from the SAW when an additional hard-shell repulsion is introduced (top right of Fig. 2[26]. The RW, SAW, and NAW all exhibit purely entropic elasticity; energetic elasticity is excluded because all configurations have identical energies. By contrast, the DJ model and ISAW introduce energy gradients via soft-core and soft-shell repulsions, respectively, with interaction energies $\varepsilon\,(>0)$ (bottom panels of Fig. 2). By varying $\varepsilon$, the DJ model (ISAW) connects the RW (SAW) at $\varepsilon=0$ to the SAW (NAW) at $\varepsilon=\infty$. These energy gradients enable the DJ model and ISAW to exhibit both energetic and entropic elasticities.
@@
where $\beta\,[\equiv 1/(k_{\mathrm{B}}T)]$ is the inverse temperature, $k_{\mathrm{B}}$ is the Boltzmann constant, $W_{n,m}(r)$ is the number of possible configurations $\omega$ for given $(n,r,m)$, and $m_{\mathrm{ub}}$ is an upper bound on $m$ (see Appendix A). The free energy, internal energy, and entropy are given by
$A(r,\beta\varepsilon)$ $=$ $-\frac{1}{\beta}\ln Z(r,\beta\varepsilon),$ (3)
$U(r,\beta\varepsilon)$ $=$ $\frac{\varepsilon}{Z(r,\beta\varepsilon)}\,\sum_{m=0}^{m_{\mathrm{ub}}}mW_{n,m}(r)\,e^{-\beta\varepsilon m},$ (4)
$S(r,\beta\varepsilon)$ $=$ $k_{B}\beta\left[A(r,\beta\varepsilon)-U(r,\beta\varepsilon)\right],$ (5)
respectively.
@@
where $\varDelta W_{n,m}(r)\equiv\left[W_{n,m}(r+\varDelta r)-W_{n,m}(r-\varDelta r)\right]/2$ and $\varDelta^{2}W_{n,m}(r)\equiv W_{n,m}(r+\varDelta r)-2W_{n,m}(r)+W_{n,m}(r-\varDelta r)$. Here, $\varDelta r\equiv 2$ because $\omega$ exists only when $r$ and $n$ are both even or both odd. The stiffness $k$ is the sum of energetic ($k_{U}$) and entropic ($k_{S}$) contributions, where $k_{U}(r,\beta\varepsilon)\equiv\partial^{2}U(r,\beta\varepsilon)/\partial r^{2}$ and $k_{S}(r,\beta\varepsilon)\equiv-T\partial^{2}S(r,\beta\varepsilon)/\partial r^{2}$. We use
$k_{S}(r,\beta\varepsilon)=-\beta\varepsilon\frac{\partial k(r,\beta\varepsilon)}{\partial\beta\varepsilon},$ (7)
which is derived from Maxwell’s relations [1, 4, 5, 10], and $k_{U}=k-k_{S}$. When it is necessary to distinguish the models, we employ “RW” or “SAW” as superscripts on $W_{n,m}(r)$, and “DJ” and “ISAW” on other physical quantities.
III Exact enumeration for the DJ model and ISAW
@@
We exactly enumerated $W^{\mathrm{RW}}_{n,m}(r)$ for $n=1,\dots,20$, $W^{\mathrm{RW}}_{n,m}(n-8)$ for $n=21,22,23$, $W^{\mathrm{RW}}_{n,m}(n-10)$ for $n=21,\dots,26$, and $W^{\mathrm{SAW}}_{n,m}(n-10)$ for $n=21,\dots,29$, employing the simplest recursive algorithm [28] with two pruning algorithms. Here, we considered the octahedral symmetry of the simple cubic lattice and the reachability of $\omega$ to an endpoint on the $x$ axis (the method is detailed in Sec. S2 in the Supplemental Material of Ref. [10]).
Comprehensive lists of $W^{\mathrm{RW}}_{n,m}(r)$ values obtained from the exact enumerations are provided in Sec. S1 of the Supplemental Material [27]. Table 1 provides an illustrative subset of the enumeration results for $W^{\mathrm{RW}}_{n,m}(r)$.
Table 1: List of $W^{\mathrm{RW}}_{n,m}(r)$ for $n=20$ and $r=8,10,12$, which is a part of Table S18 in Supplemental Material [27].
$m$
$r$
88 $10$ $12$
0 $14322531084$ $2625286352$ $227589504$
11 $43706172200$ $5625406136$ $303906776$
22 $63499792684$ $5648544020$ $198896588$
33 $55563554760$ $3369701616$ $75755328$
44 $39961462284$ $1872169836$ $33960789$
55 $25111433136$ $855424074$ $10036914$
66 $13673567535$ $359838406$ $3635132$
77 $7249781378$ $161594846$ $1310080$
88 $4078923694$ $73451960$ $466777$
99 $2069333612$ $32887440$ $204330$
$10$ $1139672875$ $15051476$ $67964$
$11$ $509421524$ $4755638$ $12658$
$12$ $296702202$ $3405656$ $14197$
$13$ $187872584$ $1719114$ $4520$
$14$ $111262038$ $916660$ $2236$
$15$ $36803762$ $135714$ 0
$16$ $26046711$ $201446$ $512$
$17$ $18453292$ $99040$ $88$
$18$ $8264852$ $28494$ 0
1919 $5808674$ $23132$ 0
2020 $2856074$ $9904$ $12$
2121 $1912686$ $6762$ 0
$22$ $1154864$ $2870$ 0
$23$ $249366$ 0 0
$24$ $217799$ $180$ 0
$25$ $255924$ $456$ 0
$26$ $141410$ $90$ 0
$27$ $46398$ 0 0
$28$ $40100$ 0 0
$29$ $4800$ 0 0
$30$ $11812$ $10$ 0
$31$ $5552$ 0 0
$32$ $3258$ 0 0
$33$ $140$ 0 0
$34$ $210$ 0 0
$36$ $434$ 0 0
$37$ $84$ 0 0
$42$ 88 0 0
To validate the exact enumeration results for $W_{n,m}^{\mathrm{RW}}(r)$ presented in Sec. S1 of the Supplemental Material [27], we confirm their consistency with those reported in Refs. [10, 29, 30, 31, 32, 33, 34, 35] using two independent approaches. First, we validate the values of $W_{n,m}^{\mathrm{RW}}(r)$ for $m=0$ using the equality of the number of $n$-step RWs without intersections $W_{n,0}^{\mathrm{RW}}(r)$ and the number of $n$-step SAWs $W_{n}^{\mathrm{SAW}}(r)$. We insert our enumerated values of $W_{n,0}^{\mathrm{RW}}(r)$ into the left side of $W_{n,0}^{\mathrm{RW}}(r)=W_{n}^{\mathrm{SAW}}(r)$, and the reference values of $W_{n}^{\mathrm{SAW}}(r)$ from Refs. [30, 32] and Table S1 of Ref. [10] into the right side; the agreement was perfect.
Second, we validate the values of $W_{n,m}^{\mathrm{RW}}(r)$ using the relation $W_{n}^{\mathrm{RW}}(r)=\sum_{m=0}^{m_{\mathrm{ub}}}W_{n,m}^{\mathrm{RW}}(r)$, where $W_{n}^{\mathrm{RW}}(r)$ is the total number of $n$-step RWs in a cubic lattice from the origin to the site $(r,0,0)$. To compute the reference values for $W_{n}^{\mathrm{RW}}(r)$, we adapt the formula given in Ref. [33] to our notation, with the endpoints anchored at the origin and $(r,0,0)$:
$W_{n}^{\mathrm{RW}}(r)=\binom{n}{\frac{n-r}{2}}\sum_{k=0}^{\frac{n-r}{2}}\binom{\frac{n-r}{2}}{k}\binom{\frac{n+r}{2}}{k}\binom{2k}{k}$. (8)
Equation (8) combines the number of possible step arrangements along each axis using binomial coefficients, which represent the number of possible step choices in each axial direction. Using Eq. (8), we computed the reference values for $W_{n}^{\mathrm{RW}}(r)$ that are listed in Table S24 of the Supplemental Material [27]. These reference values are in perfect agreement with the literature values for $s=n$ (i.e., $r=0$ for even $n$ and $r=1$ for odd $n$): with Ref. [29] for $n=2,4,\dots,12$; with Refs. [31, 34] for $n=14$ and $16$; and with Ref. [35] for $n=1,3,\dots,17$. This consistency across multiple independent sources validates our reference values for $W_{n}^{\mathrm{RW}}(r)$. For all $n$ and $r$ considered, there is perfect agreement between the reference values of $W_{n}^{\mathrm{RW}}(r)$ and the summations $\sum_{m=0}^{m_{\mathrm{ub}}}W_{n,m}^{\mathrm{RW}}(r)$ computed from our enumeration results in Tables S1 to S18 of the Supplemental Material [27]. This agreement validates the exact enumeration of $W_{n,m}^{\mathrm{RW}}(r)$.
IV Polynomials in arbitrary $n$ of $W^{\mathrm{RW}}_{n,m}(r)$ and $W^{\mathrm{SAW}}_{n,m}(r)$